A car is moving with speed 30m/sec on a circular path of radius 500m.Its speed is increasing at the rate if 2m/sec squre.what is the acceleration of the car at that moment?

Grade:11

1 Answers

Vikas TU
7 years ago
Centripital accelration = v^2/r = > 900/500 => 9/5 m/s^2
tangential acceleration = 2 m/s^2
Net acceleration => root( 4 + 81/25) = > root(181/25) = > 13.45/5 = > 2.69 m/s^2